# FollowSpot: Enhancing Wireless Communications via Movable Ceiling-Mounted Metasurfaces

### v1

@@ Abstract
This paper studies the optimal placement of ceiling-mounted metasurfaces (MTSs) to help focus the wireless signal beam onto the target receiver, as inspired by the theatre spotlight. We assume that a total of $M$ MTSs are deployed, and that there are $L$ possible positions for each MTS. The resulting signal-to-noise (SNR) maximization problem is difficult to tackle directly because of the coupling between the placement decisions of the different MTSs. Mathematically, we are faced with a nonlinear discrete optimization problem with $L^{M}$ possible solutions. A remarkable result shown in this paper is that the above challenging problem can be efficiently solved within $O(ML^{2}\log(ML))$ time. There are two key steps in developing the proposed algorithm. First, we successfully decouple the placement variables of different MTSs by introducing a continuous auxiliary variable $\mu$; the discrete primal variables are now easy to optimize when $\mu$ is held fixed, but the optimization problem of $\mu$ is nonconvex. Second, we show that the optimization of continuous $\mu$ can be recast into a discrete optimization problem with only $LM$ possible solutions, so the optimal $\mu$ can now be readily obtained. Numerical results show that the proposed algorithm can not only guarantee a global optimum but also reach the optimal solution efficiently.
Index Terms:
Metasurface placement, movable passive antennas, passive beamforming, discrete optimization, global optimum.
@@ I Introduction
Metasurface (MTS) [1, 2], a.k.a. intelligent reflecting surface [3, 4] or reconfigurable intelligent surface [5, 6, 7], is an emerging wireless device that consists of electromagnetic meta-atoms, designed to manipulate the phases and amplitudes of reflected signals to enhance wireless communications. While most existing works assume that MTSs are placed at the fixed positions, this paper proposes a new paradigm wherein the MTSs are installed onto the ceiling (e.g., of a smart factory) and can move around flexibly to track the target receiver, referred to as the wireless follow spot. If there are $M$ MTSs in total and each MTS can be moved to any of the $L$ possible positions, then the MTS placement problem boils down to a discrete optimization problem with $L^{M}$ possible solutions, while finding the optimal solution via exhaustive search over them can incur prohibitive complexity when $L$ and $M$ are large. To tackle this fundamental challenge, we show in this work that the global optimum for our considered problem can be attained efficiently within $O(ML^{2}\log(ML))$ time.
@@
Clearly, there would be much more free space if the MTS is placed in the air rather than on the ground, aside from the benefit of providing line-of-sight (LoS) channels. As such, combining the unmanned aerial vehicle (UAV) technology with the MTS has become a research hotspot in recent years. While [28, 29, 30] assume a hovering UAV with the MTS attached, many other works [31, 32, 33, 34, 35] allow the UAV to move around carrying the MTS. However, there are two main issues with the UAV-based MTS: (i) the limited battery life of UAV; (ii) the limited load capacity of UAV. To sidestep the above bottlenecks, another idea [36, 37, 38, 39, 40] is to install the MTS onto the ceiling if the transmission takes place indoors. Differing from the previous works [36, 37, 38, 39, 40] that deploy the MTSs at the fixed positions on the ceiling, this work allows the MTS positions to be adjusted over time (under the discrete position constraints), thus enabling a follow-spot effect when the target receiver is moving. Moreover, it is worth noting that the idea of manipulating channels through spatial positioning is also adopted in another frontier wireless technology called movable antenna [41, 42]. However, [41, 42] consider moving the active antennas of the transmitter, whereas the present work considers moving the passive antennas (namely the meta-atoms) of MTSs.
@@
We summarize the main results of this work in the following:
@@
$L$ number of candidate positions in each cell
@@
$\mathcal{A}_{m,\ell}$ an arc of unit circle used to optimize $\bm{X}$
@@
The rest of the paper is organized as follows. Section II describes the system model and problem formulation. Section III proposes a novel MTS-placement algorithm; its extension to the multi-receiver case is discussed in Section IV. Section V shows the numerical results. Finally, Section VI concludes the paper.
Notation: For a complex number $u$, denote by $\mathfrak{Re}\{u\}$ and $\angle u$ the real part and the angle of $u$, respectively. A summary of variables is provided in Table I for ease of reference.
@@ II System Model
Figure 2: In this example, the ceiling area is divided into multiple square cells. Within each cell, an MTS is deployed and has four candidate positions (indexed as 1, 2, 3, and 4) in this example. This model is considered later in simulations in Section V.
We consider the wireless transmission in an industrial manufacturing workshop in which a controller sends a stream of control data to an actuator of the automated production line; the multi-actuator case will be discussed later in Section IV. For the controller-to-actuator communication, this work suggests a new paradigm called FollowSpot. By this paradigm, the ceiling of the industrial manufacturing workshop is divided into multiple cells, with a phase-shifter-free MTS mounted onto each cell; unlike the conventional MTSs that rely on phase shifters, the MTSs in our case manipulate the reflected channels through their placement. Moreover, we assume that there is a discrete set of possible positions for the MTS within each cell. As such, the proposed FollowSpot aims to coordinate the positions of the multiple MTSs to focus the reflected beam on the actuator, as shown in Fig. 1.
We denote by $M$ the total number of cells. Each cell has $L$ candidate positions for its associated MTS. The set of all cells is =$\mathcal{M}=\{1,2,\ldots,M\}$ and the set of all candidate positions for cell $m$ is $\mathcal{L}_{m}=\{1,2,\ldots,L\}$. The MTS is indexed as $m$ if it is mounted onto the cell $m\in\mathcal{M}$. Assume that each MTS consists of $N$ meta-atoms (a.k.a. REs). Let $h_{0}\in\mathbb{C}$ be the direct channel from the controller to the actuator. When the MTS $m$ is placed at position $\ell\in\mathcal{L}_{m}$, let $f_{m,n,\ell}\in\mathbb{C}$ be the channel from the controller to the $n$th meta-atom of MTS $m$, and let $g_{m,n,\ell}\in\mathbb{C}$ be the channel from the $n$th meta-atom of MTS $m$ to the actuator. Thus, the reflected channel $h_{m,n,\ell}\in\mathbb{C}$ induced by the $n$th meta-atom of MTS $m$ with respect to position $\ell$ is given by
@@
In this work, we assume that the channel state information (CSI) $h_{0}$ and all $\{h_{m,\ell}\}$ are known a priori [^1] The acquisition of cascaded reflected channel for MTSs has been extensively studied in the literature [43, 44].. Note that we do not assume any particular channel models for $f_{m,n,\ell}$ and $g_{m,n,\ell}$.
Let $x_{m,\ell}$ be a binary placement variable (00 or $1$) denoting whether or not the MTS $m$ is deployed at position $\ell$. Moreover, we require $\sum_{\ell\in\mathcal{L}_{m}}x_{m,\ell}=1,\,\forall m$ to ensure that only a single MTS is mounted onto each cell. For the transmitted symbol $s$ and the background noise $z\sim\mathcal{CN}(0,\sigma^{2})$, the received signal $y\in\mathbb{C}$ at the actuator is given by
@@
where $\bm{X}=[x_{m,\ell}]$. Moreover, in the rest of the paper, we use $\bm{x}_{m}$ to denote the $m$th row vector of $\bm{X}$, i.e., $\bm{x}_{m}=[x_{m,1},x_{m,2},\ldots,x_{m,L}]$, which corresponds to the placement decision of MTS $m$. It is difficult to solve the above problem directly because of the discrete constraint and the objective nonlinearity. We remark that the exhaustive search is impractical because there are a total of $L^{M}$ possible solutions. As a striking result shown in this work, the global optimum of problem (5) can be reached within $O(ML^{2}\log(ML))$ time.
@@ III-A Problem Reformulation
=$\overset{(a)}{=}\left|\mu\left(h_{0}+\sum_{m\in\mathcal{M}}\sum_{\ell\in\mathcal{L}_{m}}x_{m,\ell}h_{m,\ell}\right)\right|$
@@
Figure 3: Rationale of the proposed efficient algorithm. (a): An example of the reflected channels when $M=1$ and $L=4$. As shown in (12), the optimal solution of $\bm{X}$ depends on the projections of these channel vectors onto the real axis after the rotation by $\mu$. (b): Continuing the example in (a), we show that the solution of $\bm{X}$ remains the same when $\mu$ stays within the same arc of the unit circle. The three transition points are represented by the hollow circles. In particular, notice that $h_{m,1}$ does not have its arc, because the projection of $h_{m,1}$ onto the real axis is never the largest no matter how the four vectors are rotated in (a). (c): Continuing the example in (a), we have 6 equations like (15) in total, and hence obtain 12 solutions $(\xi_{1},\ldots,\xi_{12})$—which partition the unit circle into 12 so-called $\xi$-arcs. The set of $\xi_{\ell}$ is a superset of the set of transition points. In this example, only $\xi_{1}$, $\xi_{4}$, and $\xi_{9}$ are transition points.
@@ III-B Global Solution in Polynomial Time
Let $\mu$ move along the unit circle. The effect of multiplying $\mu$ to $h_{m,\ell}$ can be thought of as rotating the vector of $h_{m,\ell}$ in the complex plane by an angle of $\angle\mu$ counterclockwise. Thus, for the current $\mu$, the solution in (12) boils down to first rotating all the vectors of $\{h_{m,1},\ldots,h_{m,L}\}$ simultaneously by an angle of $\angle\mu$ counterclockwise and then finding which rotated vector leads to the largest projection on the real axis, as illustrated in Fig. 3. Furthermore, note that the “winner” vector that achieves the largest projection remains the same $h_{m,\ell}$ when $\mu$ does not move too far. To be more specific, all those $\mu$ values giving the same solution in (12) would form an arc of the unit circle. In other words, the solution $\ell_{m}^{\star}$ in (12) equals $\ell^{\prime}$ if and only if the current $\mu$ lies in the arc associated with $\ell^{\prime}$, denoted by $\mathcal{A}_{m,\ell}$. We remark that some $h_{m,\ell}$ may not have its arc, i.e., $h_{m,\ell}$ can never be the “winner” vector regardless of the value of $\mu$; this subtle point is illustrated in Fig. 3. As a consequence, we conclude that there are at most $L$ arcs to consider for MTS $m$.
@@
Moreover, we observe from (11) that the optimal solution of $\bm{x}_{m}$ only depends on $\ell^{\star}_{m}$; so, equivalently, we can say that the optimal solution of $\bm{x}_{m}$ only depends on which arc $\mathcal{A}_{m,\ell}$ contains $\mu^{\star}$. Most importantly, since there are at most $L$ arcs, we can try putting $\mu$ in each of them and then pick the best one that yields the largest value of $g(\cdot)$ in (8). Thus, it just requires a linear complexity to obtain the optimal $\mu^{\star}$ and $\bm{X}^{\star}$. To summarize, we now know how to solve problem (5) efficiently when the arcs $\mathcal{A}_{m,\ell}$ are available. But how do we obtain the arcs $\mathcal{A}_{m,\ell}$?
@@
Instead of computing each $\mathcal{A}_{m,\ell}$ directly, we consider the transition points between the arcs on the unit circle, i.e., the solution of $\bm{x}_{m}$ in (11) alters whenever $\mu$ passes through a transition point, as shown in Fig. 3. Clearly, the transition point $\xi\in\mathcal{C}$ between the arcs of $h_{m,\ell}$ and $h_{m,\ell^{\prime}}$ must satisfy the following equation:
@@
We then consider equation (15) for all $m\in\mathcal{M}$ and all pairs of distinct $\ell,\ell^{\prime}\in\mathcal{L}_{m}$. As a consequence, there are $ML(L-1)/2$ such equations, with $ML(L-1)$ solutions. The resulting solutions are sorted counterclockwise (i.e., in an increasing order of their angles) as
@@
6:     Compute the midpoint $\hat{\mu}=e^{\mathrm{j}(\angle\xi_{k}+\angle\xi_{k+1})/2}$.
@@
11:Find out which $\xi$-arc yields the largest value of $f(\bm{X}^{\star})$, and output its corresponding $\bm{X}^{\star}$.
We now plot the sequence $(\xi_{\ell})^{ML(L-1)}_{\ell=1}$ on the unit circle as in Fig. 3. These points divide the unit circle into at most [^2] Note that $\xi_{1},\xi_{2},\ldots,\xi_{ML(L-1)}$ may not be distinct. $ML(L-1)$ arcs; these arcs are referred to as the $\xi$-arc. Because $\{\xi_{\ell}\}$ is a superset of the transition point set, each arc $\mathcal{A}_{m,\ell}$ must be composed of one or more $\xi$-arc. Intuitively, the decomposition of the unit circle by $\xi_{\ell}$ is finer than that by the transition points. In particular, it is easy to see that when $\mu$ moves around within the same $\xi$-arc, the solution of $\bm{x}_{m}$ remains the same, just like the aforementioned case of $\mathcal{A}_{m,\ell}$. Then we just try putting $\mu$ inside each $\xi$-arc (e.g., let $\mu$ be the midpoint of each $\xi$-arc), and find out which $\xi$-arc yields the best $\bm{X}^{\star}$—which must be the actual solution to the original problem (5). Regarding the complexity, the most time-consuming part is to sort out the sequence $(\xi)^{ML(L-1)}_{\ell=1}$, so the overall complexity equals $O(ML^{2}\log(ML^{2}))$, which further simplifies into $O(ML^{2}\log(ML))$. Algorithm 1 summarizes all the details.

### v2

# arXiv:2507.03918v2[cs.IT] 02 Mar 2026

@@ Abstract
This work focuses on the optimal placement of metasurfaces (MTSs) onto the ceiling of an industrial manufacturing workshop. In particular, we assume that a total of $M$ MTSs are deployed, and that there are $L$ possible positions for each MTS. The resulting signal-to-noise (SNR) maximization problem is difficult to tackle directly because of the coupling between the placement decisions of the different MTSs. Mathematically, we are faced with a nonlinear discrete optimization problem with $L^{M}$ possible solutions. A remarkable result shown in this paper is that the above challenging problem can be efficiently solved within $O(ML^{2}\log(ML))$ time. There are two key steps in developing the proposed algorithm. First, we successfully decouple the placement variables of different MTSs by introducing a continuous auxiliary variable $\mu$; the discrete primal variables are now easy to optimize when $\mu$ is held fixed, but the optimization problem of $\mu$ is nonconvex. Second, we show that the optimization of continuous $\mu$ can be recast into a discrete optimization problem with only $LM$ possible solutions, so the optimal $\mu$ can now be readily obtained. Numerical results show that the proposed algorithm can not only guarantee a global optimum but also reach the optimal solution efficiently.
@@ I Introduction
Metasurface (MTS) [1, 23], a.k.a. intelligent reflecting surface [46, 31] or reconfigurable intelligent surface [6, 16, 10], is an emerging wireless device that consists of electromagnetic meta-atoms, designed to manipulate the phases and amplitudes of reflected signals to enhance wireless communications. This work focuses on the optimal placement of MTSs onto the ceiling of an industrial manufacturing workshop that has been extensively studied [37, 21, 25]. This topic is motivated from three main perspectives:
  • Why MTS? MTS is a passive equipment in the sense that it does not emit any signals itself, so the power consumption can be significantly reduced as compared to the relay and base station, especially considering the large area of the workshop.
  • Why movable MTS? Because the blockages such as automated guided vehicle (AGV) are moving over time in the workshop environment, the shadowing effects and the blind spots are changing dynamically, so it is advisable to relocate MTSs accordingly.
  • Why mount MTS to ceiling? For a typical workshop, there is a large overhead space for the MTS deployment. Moreover, MTS on the ceiling can induce many more line-of-sight (LOS) reflection paths.
While most existing works assume that MTSs are placed at the fixed positions, this paper proposes a new paradigm wherein the MTSs can move around flexibly to track the target receiver, referred to as the wireless follow spot. If there are $M$ MTSs in total and each MTS can be moved to any of the $L$ possible positions, then the MTS placement problem boils down to a discrete optimization problem with $L^{M}$ possible solutions, while finding the optimal solution via exhaustive search over them can incur prohibitive complexity when $L$ and $M$ are large. To tackle this fundamental challenge, we show in this work that the global optimum for the single-user case can be attained efficiently within $O(ML^{2}\log(ML))$ time.
@@
Clearly, there would be much more free space if the MTS is placed in the air rather than on the ground, aside from the benefit of providing line-of-sight (LoS) channels. As such, combining the unmanned aerial vehicle (UAV) technology with the MTS has become a research hotspot in recent years. While [26, 40, 43] assume a hovering UAV with the MTS attached, many other works [19, 60, 17, 42, 9] allow the UAV to move around carrying the MTS. However, there are two main issues with the UAV-based MTS: (i) the limited battery life of UAV; (ii) the limited load capacity of UAV. To sidestep the above bottlenecks, another idea [18, 35, 14, 13, 51] is to install the MTS onto the ceiling if the transmission takes place indoors. Differing from the previous works [18, 35, 14, 13, 51] that deploy the MTSs at the fixed positions on the ceiling, this work allows the MTS positions to be adjusted over time (under the discrete position constraints), thus enabling a follow-spot effect when the target receiver is moving. It is worth noting that the idea of manipulating channels through spatial positioning stems from another frontier wireless technology called movable antennas [61, 28, 62, 5]; it has been shown in [27] that movable antennas can be realized efficiently by the micro-electro-mechanical systems (MEMS). The system model considered in this paper can be recognized as movable passive antennas mounted to ceiling. The present paper aims to eliminate the blind spots from the industrial manufacturing workshop, which are caused by the blockages such as the automated guided vehicle (AGV). Because the large-scale shadowing effect is our main consideration, the latency due to the mechanical movement is tolerable. When it comes to the fast-varying deep fading (e.g., caused by the multipath effect), the physical repositioning cannot keep pace with the channel variation, so it is more advisable to tune the phase shifts of MTSs to combat fading. As a result, the phase shift optimization is performed more frequently than the MTS relocation.
The main contributions of this work are three-fold:
@@
$L$ number of candidate positions in each zone
@@
$\mathcal{A}_{m,\ell}$ an sector of unit circle used to optimize $\bm{X}$
@@
The rest of the paper is organized as follows. Section II describes the system model and problem formulation. Section III proposes a novel MTS-placement algorithm; its extension to the multi-receiver case is discussed in Section IV. Section V shows the numerical results. Finally, Section VI concludes the paper. Throughout the paper, we denote by $\mathfrak{Re}\{u\}$ the real part of a complex number $u$, and $\angle u$ the phase. A summary of the main variables used in the paper is provided in Table I.
@@ II System Model
Figure 2: In this example, the ceiling is divided into multiple square zones. Each MTS can move around within its zone; each small square area of the zone represents a possible position of the MTS, and the current position of each MTS is highlighted in red. For instance, for a particular MTS $m$, its four possible positions are indexed by $\{1,2,3,4\}$, and it is currently placed at position $3$. This model is considered later in simulations in Section V.
We consider the wireless transmission in an industrial manufacturing workshop in which a controller sends a stream of control data to an actuator of the automated production line; the multi-actuator case will be discussed later in Section IV. For the controller-to-actuator communication, this work suggests a new paradigm called FollowSpot. By this paradigm, the ceiling of the industrial manufacturing workshop is divided into multiple zones, with a phase-shifter-free MTS mounted onto each zone. The controller coordinates the MTS by the wireline connections to ensure robust low-latency transmission. Moreover, the size of the zone of each MTS should be limited so as to reduce the delay of mechanic motion. Unlike the conventional MTSs that rely on phase shifters, the MTSs in our case manipulate the reflected channels through their placement. Moreover, we assume that there is a discrete set of possible positions for the MTS within each zone [^1] Indeed, the continuous degrees of freedom can bring additional performance gain, but at the cost of much more overhead. In that case, we must build a channel model that can accurately predict the CSI at an arbitrary position, which not only is time-consuming but also incurs large CSI error. In contrast, when each MTS is limited to a discrete set of candidate positions, we just need to measure CSI at these grids, so the CSI acquisition becomes much easier. Moreover, when each MTS is allowed to be continuously placed, the control message requires significantly more bits.. As such, the proposed FollowSpot aims to coordinate the positions of the multiple MTSs to focus the reflected beam on the actuator, as shown in Fig. 1.
We denote by $M$ the total number of zones. Each zone has $L$ candidate positions for its associated MTS. The set of all zones is $\mathcal{M}=\{1,2,\ldots,M\}$ and the set of all candidate positions for zone $m$ is $\mathcal{L}_{m}=\{1,2,\ldots,L\}$. The MTS is indexed as $m$ if it is mounted onto the zone $m\in\mathcal{M}$. Assume that each MTS consists of $N$ meta-atoms. Let $h_{0}\in\mathbb{C}$ be the direct channel from the controller to the actuator. When the MTS $m$ is placed at position $\ell\in\mathcal{L}_{m}$, let $f_{m,n,\ell}\in\mathbb{C}$ be the channel from the controller to the $n$th meta-atom of MTS $m$, and let $g_{m,n,\ell}\in\mathbb{C}$ be the channel from the $n$th meta-atom of MTS $m$ to the actuator. Thus, the reflected channel $h_{m,n,\ell}\in\mathbb{C}$ induced by the $n$th meta-atom of MTS $m$ with respect to position $\ell$ is given by
@@
In this work, we assume that the channel state information (CSI) $h_{0}$ and all $\{h_{m,\ell}\}$ are known a priori [^2] Notice that our method requires accurate CSI. But more importantly, since the reflection paths from the ceiling are line-of-sight (LOS)—which is the main motivation of using the ceiling-mounted MTS, it suffices to measure the channel for one meta-atom, with the channels of the rest meta-atoms recovered by the uniform planner array (UPA). Thus, the CSI measurement is scalable. Furthermore, the MTS placement aims to eliminate blind spots, so its input CSI is the large-scale fading component. Since the motion of the obstacles (e.g., AGV) is predictable in the workshop scenario, the large-scale CSI acquisition is not that difficult. In addition, to combat the small-scale fading, the MTS relocation is not enough, and it further requires the phase shift optimization based on the small-scale CSI.. Note that we do not assume any particular channel models for $f_{m,n,\ell}$ and $g_{m,n,\ell}$.
Let $x_{m,\ell}\in\{0,1\}$ be a binary placement variable denoting whether or not the MTS $m$ is deployed at position $\ell$. Moreover, we require $\sum_{\ell\in\mathcal{L}_{m}}x_{m,\ell}=1,\,\forall m$ to ensure that only a single MTS is mounted onto each zone. For the transmitted symbol $s$ and the background noise $z\sim\mathcal{CN}(0,\sigma^{2})$, the received signal $y\in\mathbb{C}$ at the actuator is given by
@@
where $\bm{X}=[x_{m,\ell}]$. Moreover, in the rest of the paper, we use $\bm{x}_{m}$ to denote the $m$th row vector of $\bm{X}$, i.e., $\bm{x}_{m}=[x_{m,1},x_{m,2},\ldots,x_{m,L}]$, which corresponds to the placement decision of MTS $m$. The above problem is difficult to tackle directly because of the discrete constraint and the coupling between distinct $x_{m,\ell}$ and $x_{m,\ell^{\prime}}$. Some previous works [15, 32] even thought such type of problem is NP-hard. We remark that the exhaustive search is impractical because there are a total of $L^{M}$ possible solutions. As a striking result shown in this work, the global optimum of problem (5) can be reached within $O(ML^{2}\log(ML))$ time.
Remark 1
We would like to clarify that the MTS placement coordination aims to eliminate the blind spots caused by the large-scale shadowing effect, so it suffices to consider a single-antenna channel. To further reap the gain of MIMO transmission, we need the refined CSI that captures the multipath effect. Thus, after the positions of MTSs have been determined, we can refine the CSI measure and then perform any existing beamforming method to enhance the MIMO transmission.
@@ III-A Problem Reformulation
$\overset{(a)}{=}\left|\mu\left(h_{0}+\sum_{m\in\mathcal{M}}\sum_{\ell\in\mathcal{L}_{m}}x_{m,\ell}h_{m,\ell}\right)\right|$
@@ III-B Global Solution in Polynomial Time
(a) MTS 1
(b) MTS 2
(c) Combined
Figure 3: Illustration of the proposed algorithm. Consider two MTSs each with 2 possible positions. (a): The optimal position of MTS 1 depends on the unit vector $\mu$; if $\mu$ lies in the yellow sector, $h_{1,2}$ leads to a larger $f(\bm{X})$ so we should place MTS 1 at position 2; if $\mu$ lies in the purple sector, $h_{1,1}$ is better so we should place MTS 1 at position 1. (b): Similarly, if $\mu$ lies in the blue sector, we should place MTS 2 at position 1; if $\mu$ lies in the green sector, we should place MTS 2 at position 2. (c): Now we combine the sectorizations of MTSs 1 and 2 to obtain a total of 4 sectors; we just try out each sector for $\mu$ to see which yields the largest $f(\bm{X})$, thereby obtaining the global optimum. Most importantly, for the general case, the total number of sectors after combination is $O(ML)$, so the proposed algorithm is scalable.
Figure 4: Illustration of the “fake” transition point. Recall that the optimal positions depend on which sector $\mu$ lies in. The sectors are formed by a set of straight-line cuts. Each cut passes through the origin and intersects with the unit circle at two positions; an intersection is referred to as a transition point if the optimal position of MTS changes when $\mu$ passes through it. (a): When $L=2$, there is one straight-line cut, and the two intersections are both transition points. (b): When $L=3$, there are three straight-line cuts and hence six intersections; however, three intersections (marked in gray) are between two sectors that lead to the same optimal position of MTS 1, so the optimal position does not change when $\mu$ passes through any of them; these gray intersections are referred to as “fake” transition points. (c): When $L=4$, there are six straight-line cuts and thus twelve intersections; observe that nine of them are fake transition points. For the general case, there are ML(L1)/2ML(L-1)/2 straight-line cuts and ML(L1)ML(L-1) intersections, but the number of fake transition points depends on the specific case.
@@
Let $\mu$ move along the unit circle. The effect of multiplying $\mu$ to $h_{m,\ell}$ can be thought of as rotating the vector of $h_{m,\ell}$ in the complex plane by an angle of $\angle\mu$ counterclockwise. Thus, for the current $\mu$, the solution in (12) boils down to first rotating all the vectors of $\{h_{m,1},\ldots,h_{m,L}\}$ simultaneously by an angle of $\angle\mu$ counterclockwise and then finding which rotated vector leads to the largest projection on the real axis. Furthermore, note that the “winner” vector that achieves the largest projection remains the same $h_{m,\ell}$ when $\mu$ does not move too far. To be more specific, all those unit vectors $\mu$ giving the same solution in (12) would lie within a sector of the unit disk on the complex plane, as shown in Fig. 3. In other words, the solution $\ell_{m}^{\star}$ in (12) equals $\ell^{\prime}$ if and only if the current $\mu$ lies in the sector associated with $\ell^{\prime}$, denoted by $\mathcal{A}_{m,\ell}$. We remark that some $h_{m,\ell}$ may not have its sector, i.e., $h_{m,\ell}$ can never be the “winner” vector regardless of the value of $\mu$; this subtle point is illustrated in Fig. 4. As a consequence, we conclude that there are at most $L$ sectors to consider for MTS $m$.
@@
Moreover, we observe from (11) that the optimal solution of $\bm{x}_{m}$ only depends on $\ell^{\star}_{m}$; so, equivalently, we can say that the optimal solution of $\bm{x}_{m}$ only depends on which sector $\mathcal{A}_{m,\ell}$ contains $\mu^{\star}$. Most importantly, since there are at most $L$ sectors, we can try putting $\mu$ in each of them and then pick the best one that yields the largest value of $g(\cdot)$ in (8). Thus, it just requires a linear complexity to obtain the optimal $\mu^{\star}$ and $\bm{X}^{\star}$. To summarize, we now know how to solve problem (5) efficiently when the sectors $\mathcal{A}_{m,\ell}$ are available. But how do we obtain the sectors $\mathcal{A}_{m,\ell}$?
@@
Instead of computing each $\mathcal{A}_{m,\ell}$ directly, we consider the transition points between the sectors on the unit circle, i.e., the solution of $\bm{x}_{m}$ in (11) alters whenever $\mu$ passes through a transition point, as shown in Fig. 3. Clearly, the transition point $\xi\in\mathcal{C}$ between the sectors of $h_{m,\ell}$ and $h_{m,\ell^{\prime}}$ must satisfy the following equation:
@@
We then consider equation (15) for all $m\in\mathcal{M}$ and all pairs of distinct $\ell,\ell^{\prime}\in\mathcal{L}_{m}$. As a consequence, there are ML(L1)/2ML(L-1)/2 such equations, with ML(L1)ML(L-1) solutions. The resulting solutions are sorted counterclockwise (i.e., in an increasing order of their angles) as
@@
6:  Put $\mu$ inside the $\xi_{k}$-sector.
@@
11:Find out which $\xi$-sector yields the largest value of $f(\bm{X}^{\star})$, and output its corresponding $\bm{X}^{\star}$.
We now plot the sequence $(\xi_{\ell})^{ML(L-1)}_{\ell=1}$ on the unit circle as in Fig. 4. These points divide the unit disk into at most [^3] Note that $\xi_{1},\xi_{2},\ldots,\xi_{ML(L-1)}$ may not be distinct. ML(L1)ML(L-1) sectors; these sectors are referred to as the $\xi$-sector. Because $\{\xi_{\ell}\}$ is a superset of the transition point set, each sector $\mathcal{A}_{m,\ell}$ must be composed of one or more $\xi$-sector. Intuitively, the decomposition of the unit disk by $\xi_{\ell}$ is finer than that by the transition points. In particular, it is easy to see that when $\mu$ moves around within the same $\xi$-sector, the solution of $\bm{x}_{m}$ remains the same, just like the aforementioned case of $\mathcal{A}_{m,\ell}$. Then we just try putting $\mu$ inside each $\xi$-sector and find out which $\xi$-sector yields the best $\bm{X}^{\star}$—which must be the actual solution to the original problem (5). Regarding the complexity, the most time-consuming part is to sort out the sequence $(\xi)^{ML(L-1)}_{\ell=1}$, so the overall complexity equals $O(ML^{2}\log(ML^{2}))$, which further simplifies into $O(ML^{2}\log(ML))$. Algorithm 1 summarizes all the details.